# Test-time regression: a unifying framework for designing sequence models with associative memory

### v1

@@ Abstract
where $\mathbf{M}_{t}$ is the matrix representation of ${m}_{t}$. This is exactly the recurrence of DeltaNet (Schlag et al., 2021, Yang et al., 2024c). As we perform stochastic gradient descent, we produce an online sequence of memories $\{\mathbf{M}_{i}\}_{i=1}^{t}$, each of which is an associative memory ${m}_{i}$ for key-value pairs up to $(\mathbf{k}_{i},\mathbf{v}_{i})$. The stochastic gradient approach to streaming inputs has the advantage that ${m}_{t}$ focuses more on recent key-value pairs, rather than uniformly over all timesteps as in the case of unweighted RLS and unweighted linear attention. This recency bias allows online learners to handle non-stationary data such as text and time-series.
Equation 18 is known by many names, such as the delta rule or the Widrow-Hoff algorithm (Widrow and Hoff, 1988). Within the adaptive filtering literature this algorithm is known as Least-Mean Squares (LMS) (Haykin, 2014); there it is one of the most common algorithms for handling non-stationary signals. Classic neurocomputing literature (Kohonen, 1989, Equation 6.33) has also previously studied using gradient descent with a single example to update an associative memory map. The iterates $\mathbf{M}_{t}$ are also known as “fast weights”, which were originally proposed by Schmidhuber (1992) as an internal learning/optimization procedure executed at test time, in contrast to the “slow weights” of a neural network that are updated only at training time. It is also possible to apply Equation 16 to iteratively produce nonlinear regressors (Sun et al., 2024).
Adaptive step sizes.
@@
Adaptive step sizes is an intuitive improvement to our gradient-descent-based recurrent architecture. A common way to derive adaptive step sizes in general is by leveraging the relationship between gradient descent and online learning (Cesa-Bianchi et al., 2004, Zinkevich, 2003), a connection that has produced many successful algorithms in the past, such as Adagrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2015). Recently, Liu et al. (2024) derived an effective sequence layer called Longhorn from an online learning perspective (Kulis and Bartlett, 2010), solving
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\ \frac{1}{2}\lVert\mathbf{M}-\mathbf{M}_{t-1}\rVert_{F}^{2}+\frac{\delta_{t}}{2}\lVert\mathbf{v}_{t}-\mathbf{M}\mathbf{k}_{t}\rVert_{2}^{2},$ (19)
resulting in the recurrence
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}\left(\mathbf{I}-\frac{1}{1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2}}{\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}}\right)+\frac{1}{1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2}}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}.$ (20)
@@
Comparing the Longhorn update in Equation 20 to the SGD update in Equation 18, we see that Longhorn’s layer is equivalent to SGD with an adaptive step size of $\beta_{t}=1/(1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2})$. This alternative perspective reflects the fact that online learning is essentially equivalent to stochastic optimization (Cesa-Bianchi et al., 2004, Duchi et al., 2011, Hazan, 2022).
In Equation 19, $\delta_{t}$ controls how well $\mathbf{M}_{t}$ memorizes the latest association $(\mathbf{k}_{t},\mathbf{v}_{t})$. When $\delta_{t}\to\infty$, Equation 19 is equivalent to a constrained optimization problem
@@
Solving this constrained optimization problem results in a step size $\beta_{t}=1/\lVert\mathbf{k}_{t}\rVert_{2}^{2}$ with update
@@
This special case is also known as the normalized LMS (NLMS) update in adaptive filtering (Castoldi and de Campos, 2009) and known as the projection method (Tanabe, 1971) or Kaczmarz’s method (Kaczmarz, 1937) in optimization theory.
Although here we discuss test-time regressors with adaptive step sizes that only depend on the current timestep, more sophisticated step sizes with long term dependencies, such as Adam (Kingma and Ba, 2015), has been explored by Clark et al. (2022) but are more challenging to parallelize. Given that adaptive optimizers are now a staple of modern deep learning, having come a long way from naive gradient descent, an interesting future line of research would be to integrate better adaptive step sizes into test-time regressors.
L2 regularization and momentum.
In addition to adaptive step sizes, we can add regularization and momentum. We focus on L2 regularization/weight decay here, but concurrent work by Behrouz et al. (2024) derives the case with momentum. Consider L2 regularizing Equation 8 to produce the objective function
$\mathcal{L}_{\text{reg},t}(\mathbf{M})=\sum_{i=1}^{t}L_{i}(\mathbf{M})+\frac{\lambda_{t}}{2}\lVert\mathbf{M}\rVert_{2}^{2}$ (23)
where $\lambda_{t}\in[0,1]$ is the regularization strength. When we minimize this regularized objective with single-example stochastic gradient descent, we obtain the recurrence
@@
Recently, Yang et al. (2024a) proposed Gated DeltaNet motivated by combining DeltaNet (Schlag et al., 2021) with the forget gating of Mamba-2 (Dao and Gu, 2024), though with a different recurrence $\mathbf{M}_{t}=\mathbf{M}_{t-1}\alpha_{t}\left(\mathbf{I}-\eta_{t}\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}\right)+\eta_{t}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}$. Here we show that the Gated DeltaNet recurrence is actually an equivalent reparameterization of our recurrence in Equation 26:
@@
Thus Gated DeltaNet is equivalent to coupling L2 regularization with the step size and rescaling $\mathbf{v}_{t}$ by $\alpha_{t}$. We can transform Equation 30 into Equation 26 via the invertible mapping $(\alpha_{t},\eta_{t},\mathbf{v}_{t}^{\prime})=(1-\beta_{t}\lambda_{t},\beta_{t}/(1-\beta_{t}\lambda_{t}),(1-\beta_{t}\lambda_{t})\mathbf{v}_{t})$. The inverse map is given by $(\lambda_{t},\beta_{t},\mathbf{v}_{t})=((1-\alpha_{t})/(\alpha_{t}\eta_{t}),\alpha_{t}\eta_{t},\allowbreak\mathbf{v}^{\prime}_{t}/\alpha_{t})$. Past works have shown that decoupling weight decay from the step size results in better generalization and easier-to-tune hyperparameters (Loshchilov and Hutter, 2019), suggesting that similar modifications to Gated DeltaNet may improve model performance.
@@
We start first from kernel regression as most common nonparametric regression method (Schölkopf and Smola, 2001), applying the kernel trick to our feature-mapped linear memory. Let $k:\mathbb{R}^{{D_{k}}\times{D_{k}}}\to\mathbb{R}$ be a Mercer kernel such that $k(\mathbf{k}_{i},\mathbf{k}_{j})=(\Phi\Phi{}^{\top})_{ij}$ where $\Phi=\begin{bmatrix}\phi(\mathbf{k}_{1})\ldots\phi(\mathbf{k}_{t})\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{\phi}}}$ and $\phi:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{\phi}}$ is an infinite-dimensional feature map. Following our recipe for model design, we make the following design choices:
@@
    Choice of weights: assign equal weight to each association
@@
    where $\phi$ is an infinite-dimensional feature map.
@@
    Choice of function class: linear functions $\mathcal{M}_{k}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\phi(\mathbf{k}),\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}\}$.
@@
The output of the optimal kernelized nonlinear associative memory induced by kernel $k$ at timestep $t$ is:
${m}_{k,\,t}(\mathbf{q})=\mathbf{V}{}^{\top}k(\mathbf{K},\mathbf{K})^{-1}k(\mathbf{K},\mathbf{q}),$ (32)
equivalent to kernel regression. Here we follow standard shorthands from the kernel learning literature: $k(\mathbf{K},\mathbf{K})$ is a shorthand for the matrix with $k(\mathbf{k}_{i},\mathbf{k}_{j})$ in entry $(i,j)$; $k(\mathbf{K},\mathbf{q})$ is a shorthand for the vector with $k(\mathbf{k}_{i},\mathbf{q})$ in entry $i$. Interestingly, when we use an exponential kernel $k(\mathbf{k}_{i},\mathbf{k}_{j})=\exp(\mathbf{k}_{i}{}^{\top}\mathbf{k}_{j}/\sqrt{{D_{k}}})$, corresponding to an infinite-dimensional feature map, and make the crude approximation $k(\mathbf{K},\mathbf{K})=\mathbf{I}$, we obtain an associative memory equivalent to an unnormalized softmax attention. However unnormalized softmax attention is empirically known to be unstable to train (Lu et al., 2021). Thus we do not stop our derivations here; instead, we will see how the standard normalized softmax attention comes from another kind of nonparametric regressor.
@@
Local polynomial regressors is another classic class of nonparametric estimators (Wasserman, 2006), which have been historically used as nonlinear smoothing functions in data analysis (Fan, 2018). Given a query $\mathbf{q}\in\mathbb{R}^{D_{k}}$ at time $t$ that we would like to use to retrieve memory/make a prediction, let our function class be $\mathcal{M}_{\mathbf{q}}^{p}$, the set of all vector-valued functions that are locally a degree $p$ polynomial around points similar to $\mathbf{q}$, as measured by some similarity function $s$ that only depends on the distance to $\mathbf{q}$, i.e. $s(\mathbf{k},\mathbf{q})=s(\mathbf{k}-\mathbf{q})$. A classic example is the exponential smoothing kernel $s_{\text{exp},\,B}(\mathbf{k},\mathbf{q})=\exp(-\lVert\mathbf{k}-\mathbf{q}\rVert_{2}^{2}/B)$ where $B$ is a bandwidth parameter controlling the local neighborhood size around $\mathbf{q}$. Typically local polynomial regression is done for 1-dimensional data; here we generalize it to the multivariate case.
@@
    ${m}_{(p,\mathbf{q}),\,t}=\operatorname*{\arg\!\min}_{{m}\in\mathcal{M}^{p}_{\mathbf{q}}}\frac{1}{2}\sum_{i=1}^{t}\gamma_{i}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2}$ (33)
    with weights given by $\gamma_{i}=s(\mathbf{k}_{i},\mathbf{q})=s(\mathbf{k}_{i}-\mathbf{q})$. We choose weights that are monotonic in distance to ensure that ${m}_{(p,\mathbf{q}),\,t}$ is an accurate approximation within a local neighborhood of $\mathbf{q}$ defined by $s$.
@@
    Choice of function class: $\mathcal{M}_{\mathbf{q}}^{p}$, the class of polynomials around $\mathbf{q}$, such that any ${m}\in\mathcal{M}_{\mathbf{q}}^{p}$ can be written as:
    ${m}(\mathbf{k}_{i})=\mathbf{M}^{(0)}+\mathbf{M}^{(1)}(\mathbf{k}_{i}-\mathbf{q})+\mathbf{M}^{(2)}(\mathbf{k}_{i}-\mathbf{q},\mathbf{k}_{i}-\mathbf{q})+\ldots+\mathbf{M}^{(p)}(\underbrace{\mathbf{k}_{i}-\mathbf{q},\ldots,\mathbf{k}_{i}-\mathbf{q}}_{p\ \text{arguments}})$ (34)
    where $\mathbf{M}^{(j)}$ is an order $j$ multilinear map [^4] See e.g. lecture notes by Conrad for the multivariate Taylor expansion case.. This can be thought of as an order $p$ Taylor expansion approximation of the true key-value mapping around $\mathbf{q}$.
@@
    Choice of minimization algorithm: analytical solution. Solving for the minimum is equivalent to solving for the set of tensors $(\mathbf{M}^{(0)},\ldots,\mathbf{M}^{(p)})$ that minimize the least squares objective. Our prediction on $\mathbf{q}$ is then ${m}_{(p,\mathbf{q}),\,t}(\mathbf{q})=\mathbf{M}^{(0)}$.
Unfortunately, solving Equation 33 is computationally expensive for large $p$. Instead, we can restrict ourselves to the simplest case of $p=0$, corresponding to the class of locally constant functions around $\mathbf{q}$, also known as Nadaraya-Watson estimators. With this simplification, the output of our nonparametric associative memory is
$m_{(0,\mathbf{q}),\,t}(\mathbf{q})$ $=\operatorname*{\arg\!\min}_{\mathbf{M}^{(0)}\in\mathbb{R}^{D_{k}}}\frac{1}{2}\sum_{i=1}^{t}s(\mathbf{k}_{i},\mathbf{q})\lVert\mathbf{v}_{i}-\mathbf{M}^{(0)}\rVert_{2}^{2}=\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{v}_{i}.$ (35)
We may be disappointed at first that $s$ should be a function of the distance $\lVert\mathbf{k}-\mathbf{q}\rVert_{2}$ to enforce the local approximation property of the estimator in Equation 33; in contrast, softmax attention relies on an exponentiated inner product. However, if we normalize keys and queries such that $\lVert\mathbf{k}\rVert_{2}=\lVert\mathbf{q}\rVert_{2}=1$, the exponential smoothing kernel with bandwidth $B=2\sqrt{{D_{k}}}$ is equivalent to the exponential scaled dot product:
@@
$\begin{bmatrix}\mathbf{M}^{(0)}&\mathbf{M}^{(1)}\end{bmatrix}$ $=(\mathbf{V}{}^{\top}\mathbf{W}\mathbf{X})(\mathbf{X}{}^{\top}\mathbf{W}\mathbf{X})^{-1},$ $\mathbf{W}$ $\triangleq\operatorname{diag}\left(s(\mathbf{k}_{1},\mathbf{q}),\ldots,s(\mathbf{k}_{t},\mathbf{q})\right)$ (37)
$\mathbf{X}$ $=\begin{bmatrix}1&\ldots&1\\ \mathbf{k}_{1}-\mathbf{q}&\ldots&\mathbf{k}_{t}-\mathbf{q}\end{bmatrix}{}^{\top}$ $\mathbf{V}$ $=\begin{bmatrix}\mathbf{v}_{1}&\ldots&\mathbf{v}_{t}\end{bmatrix}{}^{\top},$ (38)
where we recall that $\mathbf{M}^{(0)}\in\mathbb{R}^{{D_{v}}}$ and $\mathbf{M}^{(1)}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}$. An alternate form for the prediction $\mathbf{M}^{(0)}$ that is more intuitive is
$\mathbf{M}^{(0)}$ $=\bar{\mathbf{v}}-\mathbf{M}^{(1)}\bar{\mathbf{d}},$ $\bar{\mathbf{v}}$ $\triangleq\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{v}_{i},$ $\bar{\mathbf{d}}$ $\triangleq\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{d}_{i},$ $\mathbf{d}_{i}$ $\triangleq\mathbf{k}_{i}-\mathbf{q},$ (39)

### v3

@@ Abstract
As we perform stochastic gradient descent, we produce an online sequence $\{\mathbf{M}_{i}\}_{i=1}^{t}$, each of which is an associative memory ${m}_{i}$ for key-value pairs up to $(\mathbf{k}_{i},\mathbf{v}_{i})$. The stochastic gradient approach to streaming inputs has the advantage that ${m}_{t}$ focuses more on recent key-value pairs, rather than uniformly over all timesteps as in the case of unweighted RLS and unweighted linear attention.
Equation 29 is known by many names, such as the delta rule or the Widrow-Hoff algorithm (Widrow and Hoff, 1988). Within the adaptive filtering literature this recurrence is known as Least-Mean Squares (LMS) (Haykin, 2014), serving as one of the most common algorithms for streaming signals processing. From this derivation, we see that the prevalence of the update across disciplines comes from the fact that the update rule is actually an SGD update.
In our case of sequence layers, the SGD update occurs within the layer as part of the forward pass. One extension to Equation 29 is to consider SGD with nonlinear parametric regressors ${m}_{t}$. Indeed, this extension is equivalent to test-time training (TTT) layers proposed by Sun et al. (2024), which apply SGD to a nonlinear memory ${m}_{t}$, e.g. multilayer perceptrons or linear maps with layer normalizations. In contrast to our associative memory derivation here, Sun et al. (2024) motivated their method by viewing the squared loss as an self-supervised auto-encoding objective. Another extension is to take multiple gradient steps per example $(\mathbf{k}_{t},\mathbf{v}_{t})$, which was recently explored by DeltaProduct (Siems et al., 2025).
Including adaptive step sizes.
@@
To further improve the performance of a gradient-descent-based sequence layer, we can consider adaptive step sizes. A common way to derive adaptive step sizes is to use the relationship between gradient descent and online learning (Cesa-Bianchi et al., 2004, Zinkevich, 2003), a connection that has produced many successful algorithms in the past, such as Adagrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2015).
Consider the Longhorn sequence layer derived by Liu et al. (2024) by solving the online learning objective
$\mathbf{M}_{t}=\operatorname*{\arg\!\min}_{\mathbf{M}}\ \frac{1}{2}\lVert\mathbf{M}-\mathbf{M}_{t-1}\rVert_{F}^{2}+\frac{\delta_{t}}{2}\lVert\mathbf{v}_{t}-\mathbf{M}\mathbf{k}_{t}\rVert_{2}^{2}.$ (30)
This corresponds to $d(\mathbf{M}_{t-1},\mathbf{M})=\lVert\mathbf{M}-\mathbf{M}_{t-1}\rVert_{F}^{2}/2$ and $\ell_{t}(\mathbf{M})=\delta\lVert\mathbf{v}_{t}-\mathbf{M}\mathbf{k}_{t}\rVert_{2}^{2}/$ in Equation 7. The solution to Equation 30 is the recurrence
$\mathbf{M}_{t}$ $=\mathbf{M}_{t-1}\left(\mathbf{I}-\frac{\delta_{t}}{1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2}}{\mathbf{k}_{t}\mathbf{k}_{t}{}^{\top}}\right)+\frac{\delta_{t}}{1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2}}\mathbf{v}_{t}\mathbf{k}_{t}{}^{\top}.$ (31)
@@
However, by comparing the Longhorn update in Equation 31 to the SGD update in Equation 29, we see that Longhorn’s recurrence implements SGD but with a conservative adaptive step size of $\beta_{t}=\delta_{t}/(1+\delta_{t}\lVert\mathbf{k}_{t}\rVert_{2}^{2})$. Thus although the Longhorn layer was derived from an online learning perspective, its global objective is still the least squares regression objective.
We can better understand the effects of this adaptive step size by looking at $\delta_{t}$, which controls how much we care about memorizing the latest association $(\mathbf{k}_{t},\mathbf{v}_{t})$. When $\delta_{t}=0$, we see that $\mathbf{M}_{t}=\mathbf{M}_{t-1}$ and the latest association is excluded completely. When $\delta_{t}\to\infty$, Equation 30 becomes equivalent to a constrained optimization problem
@@
Solving this constrained optimization problem results in a hyperparameter-free step size $\beta_{t}=1/\lVert\mathbf{k}_{t}\rVert_{2}^{2}$ with update
@@
This special case is also known as the normalized LMS (NLMS) update in adaptive filtering (Castoldi and de Campos, 2009) and known as the projection method (Tanabe, 1971) or Kaczmarz’s method (Kaczmarz, 1937) in optimization theory. The advantage of Equation 33 is that it is hyperparameter-free, which may make it easier to tune. However, to our knowledge, the recurrence in Equation 33 has not been explored as a sequence layer.
Including L2 regularization.
In addition to adaptive step sizes, we can add regularization. Consider adding L2 regularizing to the least squares objective Equation 18
$\mathcal{L}_{\text{reg},t}(\mathbf{M})=\sum_{i=1}^{t}L_{i}(\mathbf{M})+\frac{\lambda_{i}}{2}\lVert\mathbf{M}\rVert_{2}^{2}$ (34)
where $\lambda_{i}\in[0,1]$ is the regularization strength. When we minimize this regularized objective with single-example stochastic gradient descent like in Equation 29, we obtain the recurrence
@@
Recently, Yang et al. (2024a) proposed Gated DeltaNet motivated by combining DeltaNet (Schlag et al., 2021) with the forget gating of Mamba-2 (Dao and Gu, 2024), following the online learning perspective of Liu et al. (2024). Here we show that the Gated DeltaNet recurrence is actually an equivalent reparameterization of our recurrence in Equation 37. Starting with Equation 37, we have
@@
Thus Gated DeltaNet is equivalent to L2-regularized regression via SGD where the regularization strength $\lambda_{t}$ is coupled with the step size $\beta_{t}$ and $\mathbf{v}_{t}$ is rescaled by $\alpha_{t}$. We can transform Equation 41 into Equation 37 via the invertible mapping $(\alpha_{t},\eta_{t},\mathbf{v}_{t}^{\prime})=(1-\beta_{t}\lambda_{t},\beta_{t}/(1-\beta_{t}\lambda_{t}),(1-\beta_{t}\lambda_{t})\mathbf{v}_{t})$. The inverse map is given by $(\lambda_{t},\beta_{t},\mathbf{v}_{t})=((1-\alpha_{t})/(\alpha_{t}\eta_{t}),\alpha_{t}\eta_{t},\allowbreak\mathbf{v}^{\prime}_{t}/\alpha_{t})$. Past works have shown that decoupling weight decay from the step size results in better generalization and easier-to-tune hyperparameters (Loshchilov and Hutter, 2019), suggesting that similar modifications to Gated DeltaNet may improve model performance.
@@
We start first from kernel regression as the most common nonparametric regression method (Schölkopf and Smola, 2001), and applying the kernel trick to our feature-mapped linear memory. Let $k:\mathbb{R}^{{D_{k}}\times{D_{k}}}\to\mathbb{R}$ be a Mercer kernel such that $k(\mathbf{k}_{i},\mathbf{k}_{j})=\phi(\mathbf{k}_{i}){}^{\top}\phi(\mathbf{k}_{j})=(\Phi\Phi{}^{\top})_{ij}$ where $\Phi=\begin{bmatrix}\phi(\mathbf{k}_{1})\ldots\phi(\mathbf{k}_{t})\end{bmatrix}{}^{\top}\in\mathbb{R}^{t\times{D_{\phi}}}$ and $\phi:\mathbb{R}^{D_{k}}\to\mathbb{R}^{D_{\phi}}$ is a possibly infinite-dimensional feature map. Following our recipe for model design, we make the following design choices:
@@
    Choice of weights: assign equal weight $\gamma_{i}^{(t)}=1$ to each association
@@
    where $\phi$ is possibly an infinite-dimensional feature map.
@@
    Choice of function class: linear functions $\mathcal{M}_{k}=\{{m}\mid{m}(\mathbf{k})=\mathbf{M}\phi(\mathbf{k}),\ \mathbf{M}\in\mathbb{R}^{{D_{v}}\times{D_{\phi}}}\}$.
@@
The output of the optimal kernelized nonlinear associative memory induced by the kernel $k$ at timestep $t$ is the standard kernel regression prediction equation (Schölkopf and Smola, 2001)
${m}_{k,\,t}(\mathbf{q})=\mathbf{V}_{t}{}^{\top}k(\mathbf{K}_{t},\mathbf{K}_{t})^{-1}k(\mathbf{K}_{t},\mathbf{q}).$ (43)
Here we follow standard shorthand notation from the kernel learning literature: $k(\mathbf{K}_{t},\mathbf{K}_{t})$ is a shorthand for the matrix with $k(\mathbf{k}_{i},\mathbf{k}_{j})$ in entry $(i,j)$; $k(\mathbf{K}_{t},\mathbf{q})$ is a shorthand for the vector with $k(\mathbf{k}_{i},\mathbf{q})$ in entry $i$. When we use an exponential kernel $k(\mathbf{k}_{i},\mathbf{k}_{j})=\exp(\mathbf{k}_{i}{}^{\top}\mathbf{k}_{j}/\sqrt{{D_{k}}})$, corresponding to an infinite-dimensional feature map, and make the crude approximation $k(\mathbf{K}_{t},\mathbf{K}_{t})\approx\mathbf{I}$, we obtain
${m}_{k,t}(\mathbf{q})$ $=\mathbf{V}_{t}{}^{\top}k(\mathbf{K}_{t},\mathbf{K}_{t})^{-1}k(\mathbf{K}_{t},\mathbf{q})$ (44)
$\approx\mathbf{V}_{t}{}^{\top}k(\mathbf{K}_{t},\mathbf{q})$ (45)
$=\sum_{i=1}^{t}\mathbf{v}_{i}\,k(\mathbf{k}_{i},\mathbf{q})$ (46)
$=\sum_{i=1}^{t}\mathbf{v}_{i}\exp\left(\frac{\mathbf{k}_{i}{}^{\top}\mathbf{q}}{\sqrt{{D_{k}}}}\right),$ (47)
an associative memory map equivalent to an unnormalized softmax attention. This relationship with kernel regression has been used by many previous works to derive more efficient versions of self-attention (Peng et al., 2020, Chen et al., 2021, Choromanski et al., 2020, Lu et al., 2021, Chen et al., 2024), borrowing techniques from kernel approximation methods. However unnormalized softmax attention is empirically known to be unstable to train (Lu et al., 2021). Thus we do not stop our derivations here; instead, we will see how the standard normalized softmax attention comes from another kind of nonparametric regressor.
@@
Local polynomial regression is another classic class of nonparametric estimators (Wasserman, 2006), which have been historically used as nonlinear smoothing functions in data analysis (Fan, 2018). Here, we consider the function class $\mathcal{M}^{p}$, which is the set of all vector-valued functions that are locally approximated by a degree $p$ polynomial around each point $\mathbf{q}\in\mathbb{R}^{D_{k}}$. Here locality is measured by some similarity function $s$ that only depends on the distance to $\mathbf{q}$, i.e. $s(\mathbf{k},\mathbf{q})=s(\lVert\mathbf{k}-\mathbf{q}\rVert_{2})$. A classic example is the exponential smoothing kernel $s_{\text{exp},\,B}(\mathbf{k},\mathbf{q})=\exp(-\lVert\mathbf{k}-\mathbf{q}\rVert_{2}^{2}/B)$ where $B$ is a parameter controlling the local neighborhood size around $\mathbf{q}$. Local polynomial regression is most often discussed for one-dimensional data, but here we generalize it to the multivariate case.
@@
    ${m}_{p,\,t}(\mathbf{q})=\operatorname*{\arg\!\min}_{{m}\in\mathcal{M}^{p}}\frac{1}{2}\sum_{i=1}^{t}\gamma_{i}^{(t)}\lVert\mathbf{v}_{i}-{m}(\mathbf{k}_{i})\rVert_{2}^{2}$ (48)
    with weights given by $\gamma_{i}^{(t)}=s(\mathbf{k}_{i},\mathbf{q})=s(\lVert\mathbf{k}_{i}-\mathbf{q}\rVert_{2})$.
@@
    Choice of function class: $\mathcal{M}^{p}$, the class of functions ${m}_{p}$ such that for all points $\mathbf{q}\in\mathbb{R}^{D_{k}}$, the function can be locally approximated by a polynomial
    $\hat{{m}}_{\mathbf{q}}(\mathbf{k})=\mathbf{M}^{[0]}+\mathbf{M}^{[1]}(\mathbf{k}-\mathbf{q})+\mathbf{M}^{[2]}(\mathbf{k}-\mathbf{q},\mathbf{k}-\mathbf{q})+\ldots+\mathbf{M}^{[p]}(\underbrace{\mathbf{k}-\mathbf{q},\ldots,\mathbf{k}-\mathbf{q}}_{p\ \text{arguments}})$ (49)
    where $\mathbf{M}^{[j]}$ is an order $j$ multilinear map. Intuitively, $\mathcal{M}^{p}$ is the class of functions that can be well approximated by an order $p$ Taylor expansion around each point. Consequently, the value retrieved from ${m}_{p,\,t}$ at a point $\mathbf{q}$ is given by the constant term of the polynomial: ${m}_{p}(\mathbf{q})=\hat{m}_{\mathbf{q}}(\mathbf{q})=\mathbf{M}^{[0]}$.
@@
    Choice of minimization algorithm: analytical solution. Solving for the minimum is equivalent to solving for the set of tensors $(\mathbf{M}^{[0]},\ldots,\mathbf{M}^{[p]})$ that minimize the least squares objective. Note that we need to jointly solve for all of $(\mathbf{M}^{[0]},\ldots,\mathbf{M}^{[p]})$ to get the retrieved value $\mathbf{M}^{[0]}={m}_{p,\,t}(\mathbf{q})$.
Let us consider the simplest case of $p=0$, corresponding to the class of locally constant functions around $\mathbf{q}$, also known as Nadaraya-Watson estimators (Fan, 2018, Zhang et al., 2024, Goel and Bartlett, 2024). With this simplification, the output of our nonparametric associative memory is
$m_{0,\,t}(\mathbf{q})$ $=\operatorname*{\arg\!\min}_{\mathbf{M}^{[0]}\in\mathbb{R}^{D_{k}}}\frac{1}{2}\sum_{i=1}^{t}s(\mathbf{k}_{i},\mathbf{q})\lVert\mathbf{v}_{i}-\mathbf{M}^{[0]}\rVert_{2}^{2}=\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{v}_{i}.$ (50)
We may be disappointed at first that $s$ needs to be a monotonic function of the distance $\lVert\mathbf{k}-\mathbf{q}\rVert_{2}$ to enforce the local approximation property of the estimator in Equation 48; in contrast, softmax attention relies on an exponentiated inner product. However, if we normalize keys and queries such that $\lVert\mathbf{k}\rVert_{2}=\lVert\mathbf{q}\rVert_{2}=1$, the exponential smoothing kernel with bandwidth $B=2\sqrt{{D_{k}}}$ is equivalent to the exponential scaled dot product:
@@
By deriving self-attention from the same framework as the previous recurrent sequence layers, we can more clearly compare them through the perspective of how they do associative recall. Previous sequence layers that we derived maintained an explicit matrix-valued recurrent state to compress the prefix key-value pairs, relying on this constant-sized matrix to perform retrieval. In contrast, self-attention stores the entire set of prefix key-value pairs as its associative memory system, using all of them for retrieval. As a result, its memory (and expressiveness) grows with the sequence length, a typical property of non-parametric methods (Wang et al., 2019, Wasserman, 2006).
@@
$\begin{bmatrix}\mathbf{M}^{[0]}&\mathbf{M}^{[1]}\end{bmatrix}$ $=(\mathbf{V}_{t}{}^{\top}\mathbf{W}_{t}\mathbf{X}_{t})(\mathbf{X}_{t}{}^{\top}\mathbf{W}_{t}\mathbf{X}_{t})^{-1},$ $\mathbf{W}_{t}$ $\triangleq\operatorname{diag}\left(s(\mathbf{k}_{1},\mathbf{q}),\ldots,s(\mathbf{k}_{t},\mathbf{q})\right)$ (52)
$\mathbf{X}_{t}$ $=\begin{bmatrix}1&\ldots&1\\ \mathbf{k}_{1}-\mathbf{q}&\ldots&\mathbf{k}_{t}-\mathbf{q}\end{bmatrix}{}^{\top}$ $\mathbf{V}_{t}$ $=\begin{bmatrix}\mathbf{v}_{1}&\ldots&\mathbf{v}_{t}\end{bmatrix}{}^{\top},$ (53)
where $\mathbf{W}_{t}\in\mathbb{R}^{t\times t},\ \mathbf{X}_{t}\in\mathbb{R}^{t\times(1+{D_{k}})}$$\mathbf{M}^{[0]}\in\mathbb{R}^{{D_{v}}}$, and $\mathbf{M}^{[1]}\in\mathbb{R}^{{D_{v}}\times{D_{k}}}$. These equations correspond to the standard prediction equations for local linear regression (Fan, 2018).
An alternate form for the prediction $\mathbf{M}^{[0]}$ that is more intuitive is
$\mathbf{M}^{[0]}$ $=\bar{\mathbf{v}}-\mathbf{M}^{[1]}\bar{\mathbf{d}},$ $\bar{\mathbf{v}}$ $\triangleq\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{v}_{i},$ $\bar{\mathbf{d}}$ $\triangleq\sum_{i=1}^{t}\frac{s(\mathbf{k}_{i},\mathbf{q})}{\sum_{j=1}^{t}s(\mathbf{k}_{j},\mathbf{q})}\mathbf{d}_{i},$ $\mathbf{d}_{i}$ $\triangleq\mathbf{k}_{i}-\mathbf{q},$ (54)